OpenStudy (anonymous):

translate the phrase to an algebraic expression.state what each variable represents. seven more than a number.

OpenStudy (highschoolmom2010):

n=m+7

OpenStudy (texaschic101):

That is correct....n = m + 7 n = total m = a number seven more then a number : m + 7